UT 20131031 DD Science Observing

ISA: MEdwards (LBTO)
OSA: GBechetti (LBTO)
ISp: E Solheid

Summary

High IT: 4830 LUCI Shield Temperature High IT: 4896 LUCI1 Web IO Instrument Control login problems

Opened late after high humidity. Started with LUCI in clear skies, but terrible (2") seeing, (potentially due to dome and mirror temps). Worked on MW's primary target with LUCI Br-gamma and K-s imaging, but not sure we were able to see anything useful in such poor conditions. Then LUCI temperature issues (detector above 77.05K) caused us to stop with LUCI and switch instruments to MODS.

Continued DD observation with 2 MODS targets for MW include one very high priority source. Seeing still terrible (1.5") but mostly everything behaved.

Details

01:50 -- Opening after being closed for humidity

02:00 -- While GB was collimating and pointing, I tested calibration unit and had a problem. Tried the standard fix through the webpage and received error message that unit is not configured for login. Calling Dave. We can't find immediate solutions. Filing IT.

02:05 -- Temperature on cold bench rising. As per Walter's instructions, were going to continue until we see the detector temperature rise above 77.05 K

02:15 -- Preset to MW-1; LUCI Ks imaging. IQ (GCS) = ~1.7" - 2.0" @ El = 66; IQ (on-image in Ks) = 1.4"

02:36 -- Preset to MW-1; LUCI Br-imaging

02:40 -- Preset failed. Guide star not available. Geno is checking pointing again. However, as seeing is improving a bit (1.5") I've decided to go back and get a bit more flux in Ks.

02:45 -- Prest to MW-1; LUCI Ks imaging. IQ = ~1.5" - 1.7" @ El = 63; IQ (on-image in Ks) ~ 1.1"

03:00 -- Preset was cancelled on jitter 8/9. Telescope problem, maybe MCSPU. Had to kill LUCI script.

03:05 -- Continued problems. LUCI did not recover nicely. Had to restart the telescope service from the engineer account. Still not taking the preset -- turning various systems (like rotator) back on. Rechecking pointing again after it could not find guide star

03:10 -- Preset to MW-1; LUCI Br-gamma Imaging. IQ = ~1.4 - 1.7" @ El = 60. IQ (on-image in Ks) ~ 1.2" (but variable)

03:24 -- Since INAF really doesn't have much for conditions this bad, we are continuing with this script. This is poor compared to the seeing MW was hoping for, but not much else to do in these conditions.

03:40 -- Seeing is beginning to improve significantly. GCS IQ ~ 1.2 - 1.5" @ El = 56; IQ (on-image in Ks) ~ 1.0"

03:56 -- LUCI temp rose rapidly! Stopping observations now that we are above 77.05 (actually jumped quickly to 77.1 while I was writing this!). Geno is trying to slew the rotator to see if we can get the temperature down.

04:20 -- Since all the INAF high priority targets are LUCI we made a decision to try to change the rotator position and get the temp down, however this failed. We are now giving up on LUCI and reconfiguring to MODS and turning over to INAF for a bit...

09:38 -- Preset to MW-2. MODS longslit. IQ = ~1.3-1.4 (GCS) @ El = 53

09:40 -- Alignment files r0036, r0037, r0038, r0039, r0040. Initial modsAlign offset x= -4.691, y =10.638. (Sent another very small (-0.05 = x) offset for 39 but preferred peak flux in 38 so removed it.

09:55 -- Observing MW-2. MODS longslit. Filenumbers: r0041, r0042, r0043, r0044, r0045, r0046; b0019, b0020, b0021, b0022, b0023, b0024

10:42 -- Reaquiring MW-2 with slightly different PA to adjust for refraction. IQ = ~1.5-1.7 (GCS) @ El = 51; Alignment files r0047, r0048, r0049m -- slight nudge for r0050

10:55 -- Observing MW-2. Filenumbers: r0051, b0025 Seeing unfortunately continued to degrade. Then preset cancelled. Decided it was getting late and better to move on.

11:05 -- Preset to MW-3. Problem with Azimuth Drive.

11:11 -- Preset to MW-3. IQ = ~1.3-1.5 (GCS) @ El = 56

11:20 -- Alignment files r0052, r0053, r0054: offset x- -2.542; y = 7.962;

11:30 -- Observing MW-3. Filenumbers: r0055, r0056, r0057; b0026, b0027, b0028 (B is really faint. Seeing is not good and target is setting as we come up on twilight)

12:01 -- Spectrophotometric Standard. Hiltner 600. Alignment images r0058, r0059, r0060

12:10 -- Spectrophotometric Standard. Hiltner 600. Images r0061, 0062, 0063; b0029, b0030, b0031. IQ = ~1.2 (GCS) @ El = 57

12:13 -- Astronomical Twilight

12:22 -- Closing up for night and proceeding with calibrations

12:35 -- grpixflats. r = 0064-0068; b = 0032-0036 & b = 0037-0041

12:23 -- slitflats. 1.0" r=0069-0074; b = 0042-0047; 1.2" r=0075-0080; b = 0048-0053

12:40 -- lamps r=0081-0083; b = 0054-0056

12:48 -- 8k and 3k Biases

-- MichelleEdwards - 31 Oct 2013
Topic revision: r2 - 31 Oct 2013, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback